40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 25 15:43:29 2016, ericq, Update, SUS, OSEM issues SRM_OSEMS.pngPRM_LR.png
    Reply  Thu Aug 25 20:07:35 2016, gautam, Update, SUS, OSEM issues - maybe resolved? 
    Reply  Tue Sep 6 18:21:13 2016, rana, Update, SUS, OSEM issues PRMswing.png
       Reply  Tue Sep 6 20:52:42 2016, gautam, Update, SUS, OSEM issues P9060254.JPG
          Reply  Wed Sep 7 17:23:26 2016, gautam, Update, SUS, PRM LR fixed for now IMG_3208.JPGIMG_3211.JPG
             Reply  Tue Oct 11 00:43:50 2016, ericq, Update, SUS, PRM LR problematic again PRM_LR_Problem.pngsince_fix.pnglast10.png
                Reply  Tue Oct 11 10:15:04 2016, Steve, Update, SUS, PRM LR problematic again PRMalfoiled.png
                   Reply  Tue Oct 11 13:30:49 2016, gautam, Update, SUS, PRM LR problematic again PRM_LR_decay.png
                      Reply  Wed Oct 12 13:34:28 2016, gautam, Update, SUS, PRM LR problematic again 08.png
Message ID: 12441     Entry time: Thu Aug 25 15:43:29 2016     Reply to this: 12443   12472
Author: ericq 
Type: Update 
Category: SUS 
Subject: OSEM issues 

We've seen for some time now that one of the PRM OSEM signals has been gone, and all of the SRM signals seem dark. We had tried squishing various cables to no avail.

Today I played some "musical satellite boxes," in an attempt to see if the problems are in the chambers or in the signal chains. That is, I swapped the OSEM cables from the vacuum feedthroughs between the satellite boxes, and observed what happened.

It seems clear that something is up with SRM inside the chamber. For PRM, it's not so clear...

  PRM Satellite SRM Satellite BS Satellite
PRM OSEMs LR low + small fluctuations LR low + small fluctuations LR low + small fluctuations
SRM OSEMs No signals No Signals No Signals
BS OSEMs LR low, flat OK OK

Somehow, issues with the LR channel follow both the PRM OSEMs and the PRM satellite box. frown

PRM LR first went dark on Jul 2nd, after the IFO was vented, but before we took any doors off (which happened on the 5th). I'm not sure what may have caused this.

SRM OSEMS first went dark on the evening of Jul 18, the day before ELOG 12310, when ITMY was moved in the same chamber. Maybe this ELOG was written about work the day before, but the sensors show disturbances over the course of hours. I think we need to double check the connections in chamber. 

ELOG V3.1.3-